Determine the amount of the 2018 standard deduction allowed in the following independent situations. In each case, assume the taxpayer is claimed as another person's dependent.

Amount of the
Standard Deduction
a. Bajnok, aged 19, reports income as follows: $2,800 from walking dogs. $
b. Lance, aged 18, records income as follows: $1,000 in cash dividends from a stock investment. $
c. Neeraj, aged 69, incurs income as follows: $400 interest on a bank savings account and $8,000 in wages from a part-time job. $
d. Bobbi, who is legally blind and 15 years old, reports income as follows: $2,400 in interest income from a certificate of deposit and $1,400 earned income. $
e. Abbigail, aged 60 and a widow, incurs income as follows: $12,100 from babysitting.
Answer #1

Answer:

The standard deduction as per the 2018 US rule is as below

S Case Standard deduction Remark
a Bajnok, aged 19, reports income as follows: $2,800 from walking dogs. 2800+350 =3150 Dependent Rule - Larger of 1050 or Income +350
b Lance, aged 18, records income as follows: $1,000 in cash dividends from a stock investment. 1000+350 = 1350 Dependent Rule - Larger of 1050 or Income +350
c Neeraj, aged 69, incurs income as follows: $400 interest on a bank savings account and $8,000 in wages from a part-time job. 13300 For aged above person, Standardd eduction = 12000+1300 = 13300, but income is less than 13300 limit, hence full income 8400 will be under standard deduction
d Bobbi, who is legally blind and 15 years old, reports income as follows: $2,400 in interest income from a certificate of deposit and $1,400 earned income. 2400+1400+350 = 4150 Larger of 2400+1400+350 or 1050, as per rule
e Abbigail, aged 60 and a widow, incurs income as follows: $12,100 from babysitting. 13600 For widow person Standardd eduction = 12000+1600 = 13600, but income is less than 13600 limit, hence full income 12100 will be under standard deduction